# Liouville polarization and the rigidity of their Lagrangian skeleta in dimension $4$

### v1

@@ 2. Polarizations of symplectic manifolds
2.1. Neighborhoods of symplectic curves in dimension $4$
@@
$\Theta|_{F\setminus\{0\}}=d\theta\,\mbox{ for each fiber $F$ }\quad\mbox{ and }\quad d\Theta=-\frac{c_{1}}{{\mathcal{A}}}\,\pi^{*}\tau\,\mbox{ on }{\mathcal{L}}.$
@@
The 1-form $d\theta$ is defined on the punctured fibers $F\setminus\{0\}$, and in a local trivialization of ${\mathcal{L}}$ over $U\subset\Sigma$ it is defined on ${\mathcal{L}}|_{U}\setminus U$, but not on all of ${\mathcal{L}}\setminus\Sigma$. A transition function between two local trivializations is of the form $(z,w)\mapsto(z,g(z)\;\!w)$, where $g$ is a smooth $\mathbb{C}\setminus\{0\}$-valued function, and $d\theta$ is pulled back by this map to $d\theta+\operatorname{im}\frac{dg}{g}$. We can therefore define the angular class $\{d\theta\}$ as the set of $1$-forms on ${\mathcal{L}}\setminus\Sigma$ that locally are equal to $d\theta$ up to a smooth 1-form on ${\mathcal{L}}$. The form $\Theta$ is a representative of $\{d\theta\}$. Recall that we have fixed a Hermitian metric on ${\mathcal{L}}$.
@@
2.2. Polarizations (closed case)
In [3], Biran introduced the notion of polarization for closed symplectic manifolds with rational symplectic class. This definition was later extended in [25] as follows.
@@ 2.2. Polarizations (closed case)
A polarization ${\bf\Sigma}:=\{(\Sigma_{i},\mu_{i})\}$ of a closed $4$-dimensional symplectic manifold $(M,\omega)$ is a finite collection of closed 2-dimensional symplectic submanifolds $\Sigma_{i}$ that intersect $\omega$-orthogonally, weighted by real coefficients $\mu_{i}>0$, such that
(2.4) $[{\bf\Sigma}]\,:=\,\sum_{i}\mu_{i}\;\![\Sigma_{i}]\,=\,\operatorname{PD}([\omega])\,\in\,H_{2}(M;\mathbb{R}).$
In other terms, a polarization is a weighted symplectic divisor with normal crossings Poincaré-dual to $[\omega]$ in the sense of (2.4). We write $\Sigma:=\bigcup_{i}\Sigma_{i}$. We say that the polarization is smooth if it consists of a single component. In this case, $\mu$ is determined by $\Sigma$ and $[\omega]$, and we say that $\Sigma$ has degree $d=1/\mu$.
Let $(M,\omega,{\bf\Sigma})$ be a polarized closed symplectic manifold. The symplectic form $\omega$ is exact on the complement of $\Sigma$, hence $\omega$ has primitives on $M\setminus\Sigma$. Throughout the paper, we consider only Liouville forms that satisfy a regularity assumption near $\Sigma$, that we discuss now.
@@
for $\varepsilon$ small enough, and is therefore endowed with a radial coordinate $R$ and a connection $1$-form $\Theta$, as in Section 2.1. When ${\bf\Sigma}$ consists of several pairwise $\omega$-orthogonal components $\Sigma_{i}$, we follow [25, Section 3.1] and make a neighbourhood of $\Sigma$ a symplectic plumbing of the $\varepsilon$-disc bundles over the $\Sigma_{i}$, which we still denote $\textnormal{\small SDB}_{\varepsilon}(\Sigma,M)$: a deleted neighbourhood of each $\Sigma_{i}$ can be endowed with local coordinates $(R_{i},\theta_{i})$, $d\theta_{i}$ belonging to the angular class around $\Sigma_{i}$, such that on a neighbourhood of an intersection point of $\Sigma_{i}$ and $\Sigma_{j}$ we have $\Sigma_{i}=\{R_{j}=0\}$, $\Sigma_{j}=\{R_{i}=0\}$ and $\omega=d(R_{i}d\theta_{i}+R_{j}d\theta_{j})$.
@@ A $1$-form $\lambda$ on $M\setminus\Sigma$ is tame
(i) The class of tame forms around $\Sigma$ depends on the choice of the Hermitian metrics on the ${\mathcal{L}}_{i}$, and on the identification of our neighbourhood of $\Sigma$ in $M$ with the plumbing of the $\varepsilon$-disc bundles over $\Sigma_{i}$. Given a polarization, we fix these choices, whether implicitely or after a construction. The notion of tameness will always refer to these choices.
@@
It is not hard to check that the number $a_{i}$ depends only on the component $\Sigma_{i}$ and not on the point $p$ on this component (see [26, Lemma 4.1]). We call $a_{i}$ the residue of $\lambda$ at $\Sigma_{i}$ and denote it $\textnormal{Res}\,(\lambda,\Sigma_{i})$. By [25, Lemma 3.2] plumbings of symplectic disc bundles have tame Liouville forms with arbitrary residues. When, moreover, the homological polarizing condition (2.4) holds for $\mu_{i}=-\textnormal{Res}\,(\lambda,\Sigma_{i})$, then these tame Liouville forms extend to Liouville forms on $M\setminus\Sigma$, see the proof of [26, Lemma 4.1 (iii)]. For further reference we state:
@@
Let $(M,\omega,{\bf\Sigma})$ be a polarized closed symplectic manifold with ${{\bf\Sigma}=\{(\Sigma_{i},\mu_{i})\}}$. Then
@@
Proof.  (i) has been discussed above and (ii) holds by definition of tameness. We prove (iii). Let $\lambda$ be a Liouville form tame along $\Sigma$. Near a regular point of $\Sigma_{i}$, $\lambda$ is then equal to $-\mu_{i}\>\!\Theta_{i}$ modulo a bounded one-form on $M\backslash\Sigma_{i}$, and we have seen that $\Theta_{i}$ is $\omega$-dual to $\frac{\partial}{\partial R_{i}}$, so
@@
Given a polarized closed symplectic manifold $(M,\omega,{\bf\Sigma})$ and a tame Liouville form $\lambda$ on $M\setminus\Sigma$, we denote
${\mathcal{B}}({\bf\Sigma},\lambda,M)$ ::=:= $\Sigma\cup\{p\in M\mid\mbox{$\exists$ $t(p)>0$,\; }\lim_{t\nearrow t(p)}\phi^{t}_{X_{\lambda}}(p)\in\Sigma\},$
$\textnormal{Skel}({\bf\Sigma},\lambda,M)$ ::=:= $M\setminus{\mathcal{B}}(M,{\bf\Sigma},\lambda).$
@@
By (iii) of Proposition 2.5, the set ${\mathcal{B}}({\bf\Sigma},\lambda,M)$ is open. It is the basin of attraction of $\Sigma$ under the Liouville flow of $\lambda$. The closed set $\textnormal{Skel}({\bf\Sigma},\lambda,M)$ is the maximal compact subset of $M\setminus\Sigma$ that is invariant under the Liouville flow. It is called the skeleton of $(M,\omega)$ with respect to the polarization ${\bf\Sigma}$ and the Liouville form $\lambda$. Finally, point (iii) also guarantees that $\overline{{\mathcal{B}}(X,\lambda,M)}\cap\Sigma=\overline{X}$.
@@
If $(\Sigma,\mu)$ is a smooth polarization of $(M,\omega)$, then the basin ${\mathcal{B}}({\bf\Sigma},\lambda,M)$ is symplectomorphic to the symplectic disc bundle $\textnormal{\small SDB}(\Sigma,M)=\textnormal{\small SDB}_{\mu}\bigl{(}\Sigma,\omega|_{\Sigma},c_{1}(T\Sigma^{\omega})\bigr{)}$. It therefore depends only very mildly on $M$ or $\lambda$ from a symplectic perspective.
@@
As before, endow the complex projective plane $\operatorname{\mathbb{C}P}^{2}$ with the Fubini–Study symplectic form $\omega_{\operatorname{FS}}$ that integrates to $1$ over every projective line. Every smooth algebraic curve in $\operatorname{\mathbb{C}P}^{2}$ is symplectic and (since $H_{2}(\operatorname{\mathbb{C}P}^{2})$ has rank $1$) is a polarization when weighted by the inverse of its algebraic degree. For instance:
@@
    When $\Sigma=\Sigma_{k}:=\{z_{0}^{k}+z_{1}^{k}+z_{2}^{k}=0\}\subset\operatorname{\mathbb{C}P}^{2}$, then $\Sigma$ is the vanishing locus of a holomorphic section $s_{k}$ of the Hermitian line bundle ${\mathcal{O}}(k)\to\operatorname{\mathbb{C}P}^{2}$ of curvature $-k\;\!\omega_{{\operatorname{FS}}}$. The form $-\frac{1}{k}d^{c}\log\|s_{k}\|$, defined on $\operatorname{\mathbb{C}P}^{2}\backslash\Sigma$, provides a tame Liouville form with residue $-\frac{1}{k}$. For $k$ even, its skeleton is the set $\Delta_{k}(\operatorname{\mathbb{C}P}^{2})$ described in Example 1.5, and for all $k$ the intersection of the skeleton with $B^{4}(1)=\operatorname{\mathbb{C}P}^{2}\setminus\operatorname{\mathbb{C}P}^{1}$ is $\Delta_{k}$.
For these examples in $\operatorname{\mathbb{C}P}^{2}$, notice the discrepancy between the “algebraic degree” of a curve and its degree as a polarizartion (the “symplectic degree”). These two degrees coincide when the Fubini–Study form is normalized so that the area of a line is $1$, but they differ for different normalizations. For instance, if $\operatorname{\mathbb{C}P}^{2}$ is obtained as the compactification of $B^{4}(2)$, so that the symplectic form becomes $2\,\omega_{\operatorname{FS}}$, a curve of algebraic degree $2$ is a polarization of degree $1$.
@@
    $\left(\operatorname{\mathbb{C}P}^{1}\times\operatorname{\mathbb{C}P}^{1},a\;\!\omega_{\operatorname{FS}}\oplus b\;\!\omega_{\operatorname{FS}}\right)$ is polarized by
@@
[…]
@@
    $\left(\operatorname{\mathbb{C}P}^{1}\times\operatorname{\mathbb{C}P}^{1},a\;\!\omega_{\operatorname{FS}}\oplus b\;\!\omega_{\operatorname{FS}}\right)$ is polarized by ${\bf\Sigma_{2}}:=$
@@
    $\left(\operatorname{\mathbb{C}P}^{2},\omega_{\operatorname{FS}}\right)$ is polarized by
@@
There exists no smooth polarization of $\operatorname{\mathbb{C}P}^{2}$ whose skeleton is the Clifford torus.
Proof.  Assume that the Clifford torus $L$ is the skeleton of a smooth polarization of $\operatorname{\mathbb{C}P}^{2}$ of degree $k$. By [2], the Gromov width of the complement is $c_{{\operatorname{G}}}(\operatorname{\mathbb{C}P}^{2}\setminus L)\leq\frac{1}{k}$. On the other hand, $\operatorname{\mathbb{C}P}^{2}\setminus L$ is symplectomorphic to the standard symplectic disc bundle $\operatorname{SDB}_{\frac{1}{k}}(\Sigma,\tau,k)$ over a closed surface of area $k$ with disc fibres of area $\frac{1}{k}$. By [24, Proposition 1.3] this disc bundle contains a symplectic ellipsoid $E(k,\frac{1}{k})$ and hence a ball $B^{4}(\frac{1}{k})$. It follows that $c_{{\operatorname{G}}}(\operatorname{\mathbb{C}P}^{2}\setminus L)=\frac{1}{k}$. This is in contradiction to the fact that $c_{{\operatorname{G}}}(\operatorname{\mathbb{C}P}^{2}\setminus L)=\frac{2}{3}$, proven in [4, p. 2887]. $\square$
2.3. Liouville polarizations (open case)

### v2

@@ 2. Polarisations of symplectic manifolds
2.1. Neighborhoods of symplectic curves in dimension 44
@@
[…]
@@
The 1-form $d\theta$ is defined on the punctured fibers $F\setminus\{0\}$, and in a local trivialisation of $\mathcal{L}$} over $U\subset\Sigma$ it is defined on ${\mathcal{L}}|_{U}\setminus U$, but not on all of ${\mathcal{L}}\setminus\Sigma$. A transition function between two local trivialisations is of the form $(z,w)\mapsto(z,g(z)\;\!w)$, where $g$ is a smooth $\mathbb{C}\setminus\{0\}$-valued function, and $d\theta$ is pulled back by this map to $d\theta+\operatorname{im}\frac{dg}{g}$. We can therefore define the angular class $\{d\theta\}$ as the set of $1$-forms on ${\mathcal{L}}\setminus\Sigma$ that locally are equal to $d\theta$ up to a smooth 1-form on $\mathcal{L}$}. The form $\Theta$ is a representative of $\{d\theta\}$. Recall that we have fixed a Hermitian metric on $\mathcal{L}$}.
@@
2.2. Polarisations (closed case)
In [3], Biran introduced the notion of polarisation for closed symplectic manifolds with rational symplectic class. This definition was later extended in [26] as follows.
@@ 2.2. Polarisations (closed case)
A polarisation ${\bf\Sigma}:=\{(\Sigma_{i},\mu_{i})\}$ of a closed $4$-dimensional symplectic manifold $(M,\omega)$ is a finite collection of closed 2-dimensional symplectic submanifolds $\Sigma_{i}$ that intersect $\omega$-orthogonally, weighted by real coefficients $\mu_{i}>0$, such that
(2.4) ${\bf\Sigma}]\,:=\,\sum_{i}\mu_{i}\;\![\Sigma_{i}]\,=\,\operatorname{PD}([\omega])\,\in\,H_{2}(M;\mathbb{R}).$
In other terms, a polarisation is a weighted symplectic divisor with normal crossings Poincaré-dual to $\omega]$ in the sense of (2.4). We write $\Sigma:=\bigcup_{i}\Sigma_{i}$. We say that the polarisation is smooth if it consists of a single component. In this case, $\mu$ is determined by $\Sigma$ and $\omega]$, and we say that $\Sigma$ has degree $d=1/\mu$.
Let $(M,\omega,{\bf\Sigma})$ be a polarised closed symplectic manifold. The symplectic form $\omega$ is exact on the complement of $\Sigma$, hence $\omega$ has primitives on $M\setminus\Sigma$. Throughout the paper, we consider only Liouville forms that satisfy a regularity assumption near $\Sigma$, that we discuss now.
@@
for $\varepsilon$ small enough, and is therefore endowed with a radial coordinate $R$ and a connection $1$-form $\Theta$, as in Section 2.1. When $\bf\Sigma$} consists of several pairwise $\omega$-orthogonal components $\Sigma_{i}$, we follow [26, Section 3.1] and make a neighbourhood of $\Sigma$ a symplectic plumbing of the $\varepsilon$-disc bundles over the $\Sigma_{i}$, which we still denote $\textnormal{\small SDB}_{\varepsilon}(\Sigma,M)$: a deleted neighbourhood of each $\Sigma_{i}$ can be endowed with local coordinates $(R_{i},\theta_{i})$, $d\theta_{i}$ belonging to the angular class around $\Sigma_{i}$, such that on a neighbourhood of an intersection point of $\Sigma_{i}$ and $\Sigma_{j}$ we have $\Sigma_{i}=\{R_{j}=0\}$, $\Sigma_{j}=\{R_{i}=0\}$ and $\omega=d(R_{i}d\theta_{i}+R_{j}d\theta_{j})$.
@@ A $1$-form $\lambda$ on $M\setminus\Sigma$ is tame
(i) The class of tame forms around $\Sigma$ depends on the choice of the Hermitian metrics on the ${\mathcal{L}}_{i}$, and on the identification of our neighbourhood of $\Sigma$ in $M$ with the plumbing of the $\varepsilon$-disc bundles over $\Sigma_{i}$. Given a polarisation, we fix these choices, whether implicitely or after a construction. The notion of tameness will always refer to these choices.
@@
It is not hard to check that the number $a_{i}$ depends only on the component $\Sigma_{i}$ and not on the point $p$ on this component (see [27, Lemma 4.1]). We call $a_{i}$ the residue of $\lambda$ at $\Sigma_{i}$ and denote it $\textnormal{Res}\,(\lambda,\Sigma_{i})$. By [26, Lemma 3.2] plumbings of symplectic disc bundles have tame Liouville forms with arbitrary residues. When, moreover, the homological polarizing condition (2.4) holds for $\mu_{i}=-\textnormal{Res}\,(\lambda,\Sigma_{i})$, then these tame Liouville forms extend to Liouville forms on $M\setminus\Sigma$, see the proof of [27, Lemma 4.1 (iii)]. For further reference we state:
@@
Let $(M,\omega,{\bf\Sigma})$ be a polarised closed symplectic manifold with ${\bf\Sigma}=\{(\Sigma_{i},\mu_{i})\}$}. Then
@@
Proof.  (i) has been discussed above and (ii) holds by definition of tameness. We prove (iii). Let $\lambda$ be a Liouville form tame along $\Sigma$. Near a regular point of $\Sigma_{i}$, $\lambda$ is then equal to $\mu_{i}\>\!\Theta_{i}$ modulo a bounded one-form on $M\backslash\Sigma_{i}$, and we have seen that $\Theta_{i}$ is $\omega$-dual to $\frac{\partial}{\partial R_{i}}$, so
@@
Given a polarised closed symplectic manifold $(M,\omega,{\bf\Sigma})$ and a tame Liouville form $\lambda$ on $M\setminus\Sigma$, we denote
${\mathcal{B}}({\bf\Sigma},\lambda,M)$ $:=$ $\Sigma\cup\{p\in M\mid\mbox{$\exists$ $t(p)>0$,\; }\lim_{t\nearrow t(p)}\phi^{t}_{X_{\lambda}}(p)\in\Sigma\},$
$\textnormal{Skel}({\bf\Sigma},\lambda,M)$ $:=$ $M\setminus{\mathcal{B}}(M,{\bf\Sigma},\lambda).$
@@
By (iii) of Proposition 2.5, the set ${\mathcal{B}}({\bf\Sigma},\lambda,M)$ is open. It is the basin of attraction of $\Sigma$ under the Liouville flow of $\lambda$. The closed set $\textnormal{Skel}({\bf\Sigma},\lambda,M)$ is the maximal compact subset of $M\setminus\Sigma$ that is invariant under the Liouville flow. It is called the skeleton of $(M,\omega)$ with respect to the polarisation $\bf\Sigma$} and the Liouville form $\lambda$. Finally, point (iii) also guarantees that $\overline{{\mathcal{B}}(X,\lambda,M)}\cap\Sigma=\overline{X}$.
@@
If $(\Sigma,\mu)$ is a smooth polarisation of $(M,\omega)$, then the basin ${\mathcal{B}}({\bf\Sigma},\lambda,M)$ is symplectomorphic to the symplectic disc bundle $\textnormal{\small SDB}(\Sigma,M)=\textnormal{\small SDB}_{\mu}\bigl(\Sigma,\omega|_{\Sigma},c_{1}(T\Sigma^{\omega})\bigr)$. It therefore depends only very mildly on $M$ or $\lambda$ from a symplectic perspective.
@@
As before, endow the complex projective plane $\operatorname{\mathbb{C}P}^{2}$ with the Fubini–Study symplectic form $\omega_{\operatorname{FS}}$ that integrates to $1$ over every projective line. Every smooth algebraic curve in $\operatorname{\mathbb{C}P}^{2}$ is symplectic and (since $H_{2}(\operatorname{\mathbb{C}P}^{2})$ has rank $1$) is a polarisation when weighted by the inverse of its algebraic degree. For instance:
@@
    When $\Sigma=\Sigma_{k}:=\{z_{0}^{k}+z_{1}^{k}+z_{2}^{k}=0\}\subset\operatorname{\mathbb{C}P}^{2}$, then $\Sigma$ is the vanishing locus of a holomorphic section $s_{k}$ of the Hermitian line bundle ${\mathcal{O}}(k)\to\operatorname{\mathbb{C}P}^{2}$ of curvature $k\;\!\omega_{{\operatorname{FS}}}$. The form $\frac{1}{k}d^{c}\log\|s_{k}\|$, defined on $\operatorname{\mathbb{C}P}^{2}\backslash\Sigma$, provides a tame Liouville form with residue $\frac{1}{k}$. For $k$ even, its skeleton is the set $\Delta_{k}(\operatorname{\mathbb{C}P}^{2})$ described in Example 1.5, and for all $k$ the intersection of the skeleton with $B^{4}(1)=\operatorname{\mathbb{C}P}^{2}\setminus\operatorname{\mathbb{C}P}^{1}$ is $\Delta_{k}$.
For these examples in $\operatorname{\mathbb{C}P}^{2}$, notice the discrepancy between the “algebraic degree” of a curve and its degree as a polarisation (the “symplectic degree”). These two degrees coincide when the Fubini–Study form is normalised so that the area of a line is $1$, but they differ for different normalisations. For instance, if $\operatorname{\mathbb{C}P}^{2}$ is obtained as the compactification of $B^{4}(2)$, so that the symplectic form becomes $2\,\omega_{\operatorname{FS}}$, a curve of algebraic degree $2$ is a polarisation of degree $1$.
@@
    $\left(\operatorname{\mathbb{C}P}^{1}\times\operatorname{\mathbb{C}P}^{1},a\;\!\omega_{\operatorname{FS}}\oplus b\;\!\omega_{\operatorname{FS}}\right)$ is polarised by
@@
    For $\lambda=(R_{1}-a)\,d\theta_{1}+(R_{2}-b)\,d\theta_{2}$ the skeleton is the single point $\{(\infty,\infty)\}$.
@@
    $\left(\operatorname{\mathbb{C}P}^{1}\times\operatorname{\mathbb{C}P}^{1},a\;\!\omega_{\operatorname{FS}}\oplus b\;\!\omega_{\operatorname{FS}}\right)$ is polarised by $\bf\Sigma_{2}$}:=
@@
    $\left(\operatorname{\mathbb{C}P}^{2},\omega_{\operatorname{FS}}\right)$ is polarised by
@@
There exists no smooth polarisation of $\operatorname{\mathbb{C}P}^{2}$ whose skeleton is the Clifford torus.
Proof.  Assume that the Clifford torus $L$ is the skeleton of a smooth polarisation of $\operatorname{\mathbb{C}P}^{2}$ of degree $k$. By [2], the Gromov width of the complement is $c_{{\operatorname{G}}}(\operatorname{\mathbb{C}P}^{2}\setminus L)\leq\frac{1}{k}$. On the other hand, $\operatorname{\mathbb{C}P}^{2}\setminus L$ is symplectomorphic to the standard symplectic disc bundle $\operatorname{SDB}_{\frac{1}{k}}(\Sigma,\tau,k)$ over a closed surface of area $k$ with disc fibres of area $\frac{1}{k}$. By [25, Proposition 1.3] this disc bundle contains a symplectic ellipsoid $E(k,\frac{1}{k})$ and hence a ball $B^{4}(\frac{1}{k})$. It follows that $c_{{\operatorname{G}}}(\operatorname{\mathbb{C}P}^{2}\setminus L)=\frac{1}{k}$. This is in contradiction to the fact that $c_{{\operatorname{G}}}(\operatorname{\mathbb{C}P}^{2}\setminus L)=\frac{2}{3}$, proven in [4, p. 2887]. $\square$
2.3. Liouville polarisations (open case)